OpenStudy (anonymous):

How much money would I have to earn in order to have $1000 net, if my gross income is taxed at 30% ? What is the mathmatical formula?

OpenStudy (anonymous):

Your net is your gross minus 30% of your gross. In numerical terms, that's: \[net=gross-\frac{30}{100}gross\]What does that equal?

OpenStudy (anonymous):

could you show me that equation solved using $1000 as the only given variable?

OpenStudy (anonymous):

at 30%

OpenStudy (anonymous):

I can't give you the whole solution, but I can help you get there. Using what I wrote above, what does the net equal?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

I need more help

OpenStudy (anonymous):

Alrighty, let's pretend your gross is 100. That would make the formula: \[net=100-\frac{30}{100}\times100\] What does that equal?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

net = 70

OpenStudy (anonymous):

Exectly! What percentage of the gross does that make the net? By which I mean, what is the percentage form of \[\frac{70}{100}\]?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

70% ?

OpenStudy (anonymous):

Yep! So the net therefore is 70% of the gross. In numerical terms, \[net=70\% \times gross\]\[net=\frac{70}{100}\times gross\] Back to the original question: If your net is $1000, what is the gross? Well, plug it into this formula: \[1000=\frac{70}{100}\times gross\]

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

am I solving for X and gross? or does the X mean (Times) the gross?

OpenStudy (anonymous):

The X means times, sorry :)

OpenStudy (anonymous):

Gross = 700

OpenStudy (anonymous):

I'm afraid not. What steps did you take to calculate that?

OpenStudy (anonymous):

Gross = 0.07

OpenStudy (anonymous):

Alas, no. Could you post your work?

OpenStudy (anonymous):

i'll try

OpenStudy (anonymous):

how do you make the 70 over 100 look like you do?

OpenStudy (anonymous):

It doesn't matter too much. Consider it magic for now. If you have \[1000=\frac{70}{100}\times gross\] Your first step would be to get rid of that fraction. Firstly, get rid of the denominator. Multiply both sides of the equation by 100: \[(100)(1000)=(100)\left(\frac{70}{100}\right) \times gross\] The 100 on the top and bottom of that fraction will cancel out, leaving you with \[100000=70 \times gross\] Next, get rid of that 70. Divide both sides by 70: \[\frac{100000}{70}=\frac{70}{70}\times gross\] Again, the 70s in that fraction on the right will cancel each other out, so you are left with \[\frac{100000}{70}=gross\]

OpenStudy (anonymous):

So the answer is 1,428.6 ?

OpenStudy (anonymous):

Yeps.

OpenStudy (anonymous):

You have helped me a great deal. I really appreciate it, and believe it or not I do not think it was magic. It really does make sense. Thanks again. Oh. and thanks for the medal.

OpenStudy (anonymous):

Good luck with your future efforts!

OpenStudy (anonymous):

Thank you